I need some help on unit multipliers.... the question is 30cm ( )( )= kilo
use two unit multipliers to convert to kilometers

Answer by stanbon(75887)

30cm (1m /100cm)(1km/1000m) = (30/10^5) = (3/10^4) = 0.0003 km
=====================
Cheers,
Stan H.